# Rolling the DICE on Idiomaticity: How LLMs Fail to Grasp Context

### v1

@@ 4.2 Results
Table 2 presents the results of model performances on our evaluation set. In it important to note that, GPT-4 was used to generate the sentences for DICE, so we do not consider its performance to be revealing.
Accuracy Comparison
With the exception of the Flan models (small to XL), which generally struggle with recognizing the idiomatic use of expressions, the other models demonstrate better performance in identifying figurative rather than literal usage. Notably, the GPT models and the LLama 3.1 (405B) model exhibit comparable results across both figurative and literal subsets, standing out as exceptions to this trend. A similar pattern is observed in both zero-shot and few-shot settings, although the one-shot setting consistently yields higher overall scores.
FLAN-XXL shows reasonable zero-shot performance but unexpectedly drops in few-shot settings, unlike GPT and LLama, which improve with more examples. We hypothesize that this decline is due to Flan models being heavily instruction-tuned and may require more specific prompt structuring in few-shot scenarios, such as clearer task-example separation or tailored formatting, compared to more advanced models like GPT and LLaMA that handle generalized prompting better.
Lenient Consistency
@@
The general trend for this evaluation metric aligns with our previous observations: models show a preference for figurative interpretations when encountering an idiom, as there is a higher proportion of idioms that the models can consistently predict to be figurative across all contextual sentences than in the literal setting. As expected, all models achieve lower scores on both subsets when evaluated based on consistency, where the model must correctly classify all variations of the same expression in each sense to be rewarded. We observe the largest aggregate drop for GPT-4o when scores are evaluated using this metric. This decline indicates that GPT-4o’s high performance (evidenced by a overall accuracy of 84.33 ± 4.44 in zershot evaluations) stems from its consistency across a broad range of idioms. However, the model lacks a deep understanding of these idioms, making it susceptible to variations. This is illustrated by an overall Lenient Consistency score of 70.32 ± 7.1, showing that the model can only accurately interpret a subset of idioms consistently across different texts. Llama 3.1 is the model with the least performance difference across the two subsets, indicating a more balanced understanding of both figurative and literal contexts.
Strict Consistency
Strict Consistency, as previously defined, requires models to correctly classify all the figurative and literal uses of an expression to be rewarded. The results from this evaluation are striking: only three models—Llama 3.1, Llama 3, and GPT4o—achieve an accuracy above 40%, with 60.36%, 57.55%, and 48.59% respectively. This indicates that while state-of-the-art models may show high performance on existing idiomaticity benchmarks, they perform very poorly when a more systematic approach is used to evaluate their understanding. Even GPT-4, which served as the annotator model, can consistently classify only 59.62% of the expressions correctly in both literal and figurative contexts. This highlights a significant gap in the current models’ ability to truly understand idiomatic expressions.
@@ 5 Impact of Frequency and Sentence Likelihood on Model Performance
This section examines how idiom frequency and sentence likelihood impact LLM performance. In human processing, idiom frequency and familiarity, alongside context, influence comprehension. Similarly, in LLMs, expression frequency and sentence probability may shape idiom detection. Exploring these factors helps clarify how language-intrinsic and model-intrinsic features affect performance.
5.1 Frequency Estimation
We use the English Web Corpus (enTenTen) (Jakubíček et al., 2013) to approximate the frequency of idioms in our dataset. This was for two key reasons: (1) it is parsed and tagged which allows us to query all morphological forms of the expressions using lemmas, ensuring comprehensive frequency counts, and (2) its large scale—52 billion words across diverse genres offers a robust generalization of natural language usage. While enTenTen is not identical to the pretraining datasets of LLMs, it provides a reasonable proxy for estimating expression frequency, as high-frequency terms in enTenTen are likely to appear frequently in similar web-based pretraining corpora. This analysis helps us understand whether model performance is affected by how often they may have encountered each expression during training.
@@ 5.2 Likelihood Scores
LLMs make predictions using likelihood estimates learned during training, which are designed to maximize the probability of their extensive training data. Recent studies have shown that models tend to perform better, or assign higher evaluation scores, to sentences they deem more likely, regardless of the specific target task or evaluation criteria (Ohi et al., 2024; McCoy et al., 2024). This is likely due to sentences with higher likelihood being encountered more frequently during training. This can affect the model’s performance, as it is more familiar with some sentences than others. This in turn highlights the limitations of the models when dealing with less common contexts.
@@ 5.3 Results
We present the frequency results for GPT-3.5 Turbo, Llama 2 (70B) and Flan-T5 XL in Figure 2. For the likelihood analysis, we present the results for Llama 3 (8B) and Flan-T5 XXL in Figure 4. Results for the other models can be found in Appendix F. These models were chosen as they are representative of the general patterns observed across all the models evaluated.
Frequency is not a free lunch!
In general, we observed a consistent pattern that accuracy correlates with expression frequency, suggesting that while models may have encountered certain expressions frequently during pretraining, they may still struggle to properly interpret their idiomaticity in new contexts. Notably, this trend was observed in 9 out of the 13 models we analyzed. We also observe that, as the frequency of an expression increases, the models tend to perform better at identifying its literal occurrences, but their accuracy in recognizing idiomatic uses declines. This could potentially be due to the literal meaning being the dominant occurrence of high-frequency expressions in the pretraining data; however, this hypothesis requires further investigation.
Sentence Likelihood
In the case of the Llama models, performance correlates positively with sentence probability. In both settings, the model perform better on sentences on which it has a higher likelihood. This is particularly the case on the literal subsets, as indicated by Llama 2 (13B) and Llama 3 (8b). For the Flan-T5 models, we see a negative or negligible correlation between frequency/likelihood and performance. As seen in §4.2, Small and Large does not appear to have effectively utilized the context to learn meanings as successfully as the other models. The counter-intuitive pattern observed in XL and XXL models could be due to models being over-confident in their wrong prediction. This leads to situations where model would assign high probabilities to idiomatic sentences but performs poorly on the idiomaticity detection task.
@@ 6 Conclusion
In this work, we contribute to idiomaticity detection in NLP with several key findings. First, we introduce DICE, a challenging dataset for context-dependent idiom detection, where distinguishing figurative from literal meanings relies heavily on understanding context. Second, we propose an evaluation framework that measures both overall accuracy and strict consistency, requiring models to correctly identify all figurative and literal instances of an expression across different contexts. Third, we show that current LLMs struggle to use context effectively, highlighting the need for models that better capture contextual nuances. Lastly, we find that while expression frequency aids recall, it is insufficient for idiomaticity detection without context, in both human cognition and model performance. We also observe a positive correlation between a model’s likelihood score and its performance on DICE, suggesting training exposure influences results, emphasizing the importance of context comprehension in LLMs.
@@ 7 Limitations
One of the limitations of our work is that some idiomatic expressions are noticeably more reliant on the context than others. This means that there were cases, where we could not provide a literal counterpart to the figurative interpretation. For example, the expression "set eyes on" has such a dominant meaning of "to see", that the annotators believed to be impossible to override. In these cases, we would discard the expression. As a result, our dataset only contains a selected sample of idioms, and we acknowledge that this idea of contrastive evaluation cannot necessarily be applied to all idioms in a language.
Another of the limitations of our work is that we only consider English idioms. We would like to have extended this work to other languages, however, this relies on the existence of idiomaticity datasets in the target languages. Moreover, the idea of making idioms literal might not be translatable to other languages, where the expression takes rigid and fixed forms.
8 Ethical Considerations
We adhere to ethical practices of data collection. All participants were required to sign a consent form and informed that they could withdraw from participation at any time without facing any consequences. Our collection procedures and processes are monitored and reviewed by a University-wide ethics committee. The committee members are unrelated and detached from this work.
@@ Acknowledgements
We acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing. This work was supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1], UK EPSRC grantEP/T02450X/1, the Turing Institute Fellowship and the UniDive COST Action. Additional thanks to our annotators for their hard work. A further acknowledgement to Lily Zeng, Thomas Pickard and members of the MWE Group for their helpful discussions.
@@ Appendix A Examples of Expert Annotations
Idiom Definition of the Figurative Meaning Sentence Accept Reject Reason (if reject)
smoking gun "a piece of incontrovertible evidence" The detective found a smoking gun at the crime scene. N Y Ambiguous
guilt trip "to make someone feel guilty" After breaking her mother’s vase, Sarah’s sister put her on a guilt trip for weeks. N Y Doesn’t make sense
turn a blind eye "pretend not to notice" Despite the obvious safety hazards, the supervisor chose to turn a blind eye. N Y Figurative
down the wire "a situation whose outcome is not decided until the very last minute" The electrician was careful not to cut down to the wire while he was working. N Y Form changed
set eyes on "see" As soon as she set eyes on the beach, she was overwhelmed by its serene beauty. N Y Skip
blow off steam "get rid of pent-up energy or emotion" During the train ride, the kids were excited to see the old locomotive blow off steam. Y
get a grip "begin to deal with or understand" He struggled to get a grip on the slippery glass jar of pickles. Y
@@ Appendix D Participant Briefing
Appendix E Implementation Details
E.1 Zero-shot Prompts
We ran the FLAN-T5 models on a NVIDIA H100 GPU. We use OpenAI’s API for interactions with the GPT models, HuggingFace for Flan-T5 models and Replicate for the Llama models. Each model was evaluated with three different prompts. All of the results we report are the average across the three prompt settings.
Model: GPT-4, GPT-4o, GPT-3.5-Turbo
Prompt 1: "Is the expression ’idiom’ used figuratively or literally in the sentence: ’sentence’. Answer ’i’ for figurative, ’l’ for literal."
Prompt 2: "In the sentence ’sentence’, is the expression ’idiom’ being used figuratively or literally? Respond with ’i’ for figurative and ’l’ for literal."
Prompt 3: "How is the expression ’idiom’ used in this context: ’sentence’. Output ’i’ if the expression holds figurative meaning, output ’l’ if the expression holds literal meaning."
Models: Flan-T5-XXL, Flan-T5-XL, Flan-T5-Large, Flan-T5-Small
Prompt 1: "Is the meaning of expression idiomatic or literal? If used idiomatically, answer ’i’, if literally, answer ’l’." "Expression: idiom" "Sentence: sentence".
Prompt 2: "In the sentence ’sentence’, is the expression ’idiom’ being used figuratively or literally? Respond with ’i’ for figurative and ’l’ for literal."
Prompt 3: "How is the expression ’idiom’ used in this context: ’sentence’. Output ’i’ if the expression holds figurative meaning, output ’l’ if the expression holds literal meaning."
Models: Llama Models
Prompt 1: Is the expression ’idiom’ used figuratively or literally in the sentence: ’sentence’. Answer ’i’ for figurative, ’l’ for literal. Put your answer after ’output: ’
Prompt 2: In the sentence ’sentence’, is the expression ’idiom’ being used figuratively or literally? Respond with ’i’ for figurative and ’l’ for literal. Put your answer after ’output: ’
Prompt 3: How is the expression ’idiom’ used in this context: ’sentence’. Output ’i’ if the expression holds figurative meaning, output ’l’ if the expression holds literal meaning. Put your answer after ’output: ’
E.2 Fewshot Prompt
We used the same examples across all model evaluations. "oneshot_example" is appended to the end of the above detailed prompts. oneshot_example: "Here is an example: The expression ’play with fire’ is used figuratively in the sentence: ’The war took away the unfortunate necessity, as Unionists saw it, to play with fire in the national interest, but it did not materially alter their view of themselves.’ → Output: i The expression ’play with fire’ is used literally in the sentence: ’Despite the danger, he decided to play with fire, poking the embers with a stick.’ → Output: l"
@@ E.3 Hyper-parameters
Table 5: Generated by Spread-LaTeX
@@
Appendix F Additional Results for Frequency and Likelihood Analysis

### v2

@@ 4.2 Results
Table 2 presents the results of model performances on our evaluation set. It is important to note that, GPT-4 was used to generate the sentences for DICE, so we do not consider its performance to be revealing.
Degradation across evaluation levels.
As expected, performance declines from Accuracy to Lenient Consistency, and further to Strict Consistency. The results from the strictest evaluation show that, only three models—Llama 3.1, Llama 3, and GPT4o—achieve an accuracy above 40%, with 60.36%, 57.55%, and 48.59% respectively. This pattern highlights that, while the current LLMs can correctly classify individual instances, they often fail to do so consistently for both literal and figurative uses of the same idiomatic expression. Compounding this, we observe substantial standard deviations across prompts, particularly for smaller models. If the models truly relied on contextual understanding, we would expect them to perform consistently across varying levels of consistency and remain robust to prompt variations; however, the results suggest that the performance is highly inconsistent, and thus, models are not effectively leveraging context.
Preference towards figurative.
The general trend for based on lenient consistency aligns with our observation on base accuracy: models show a preference for figurative interpretations when encountering an idiom, as there is a higher proportion of idioms that the models can consistently predict to be figurative across all contextual sentences than in the literal setting.
@@
We observe the largest aggregate drop for GPT-4o which indicates that GPT-4o’s high performance (evidenced by a overall accuracy of 84.33 ± 4.44 in zershot evaluations) stems from its consistency across a broad range of idioms. However, the model lacks a deep understanding of these idioms, and it is susceptible to variations. This is illustrated by an overall Lenient Consistency score of 70.32 ± 7.1, which indicates that the model can only accurately interpret a subset of idioms consistently across different texts. Llama 3.1 is the model with the least performance difference across the two subsets, indicating a more balanced understanding of both figurative and literal contexts.
Mixed impact of one-shot prompting.
Introducing an example improves performance for some models—most notably GPT-4o, which shows gains across all metrics and achieves the highest strict consistency (63.5%). However, this improvement is not consistent: for many models, particularly the Flan-T5 variants, one-shot prompting offers little benefit or even degrades performance. This likely stems from a task-specific issue: instead of clarifying the task requirements, showing examples where idiomatic expressions appear literally clashes with the model’s prior expectations. Consequently, the one-shot example creates uncertainty instead of enhancing contextual understanding.
Summary.
While larger models like GPT-4o and Llama 3.1 (405B) show stronger overall performance, the results reveal that idiomaticity disambiguation, especially under consistency-based evaluation, remains a challenging task. The substantial drop from accuracy to strict consistency, coupled with high variation across prompts, suggests that models still rely heavily on shallow heuristics rather than robust contextual understanding.
@@ 5 Impact of Frequency and Sentence Likelihood on Model Performance
This section examines how idiom frequency and sentence likelihood impact LLM performance. In human processing, idiom frequency and familiarity, alongside context, influence comprehension (Cronk et al., 1993; Levorato and Cacciari, 1992; Schweigert, 1986; Brysbaert et al., 2018). Similarly, in LLMs, expression frequency and sentence probability may shape idiom detection. Exploring these factors helps clarify how language-intrinsic and model-intrinsic features affect performance.
5.1 Frequency Estimation
We use the English Web Corpus (enTenTen) (Jakubíček et al., 2013) to approximate the frequency of idioms in our dataset. This was for two key reasons: (1) it is parsed and tagged which allows us to query all morphological forms of the expressions using lemmas, ensuring comprehensive frequency counts, and (2) its large scale 52 billion words across diverse genres offers a robust generalization of natural language usage. While enTenTen is not identical to the pretraining datasets of LLMs, it provides a reasonable proxy for estimating expression frequency, as high-frequency terms in enTenTen are likely to appear frequently in similar web-based pretraining corpora. This analysis helps us understand whether model performance is affected by how often they may have encountered each expression during training. See Appendix F for further details on how we obtain frequency counts.
@@ 5.2 Likelihood Scores
LLMs are trained to maximize the likelihood of their training data, learning to assign higher probabilities to sequences that resemble those seen during training. At inference time, these models can assign likelihood scores to input sentences, reflecting how typical or expected a sentence appears under the model’s internal distribution. Recent studies have shown that models tend to perform better, or assign higher evaluation scores, to sentences they deem more likely, regardless of the specific target task or evaluation criteria (Ohi et al., 2024; McCoy et al., 2024). In our setting, we use sentence-level likelihood as a proxy to test whether model performance on idiomaticity disambiguation is influenced by how probable a sentence appears to the model. Specifically, we ask whether models are more accurate on DICE when the input context has high likelihood, potentially indicating a preference for familiar or prototypical constructions over true contextual reasoning.
To analyze the relationship between sentence likelihood and model performance, we compute the likelihood of each sentence in DICE using the standard language modeling formulation. For a sentence $y=[y^{(1)},\dots,y^{(T)}]$, the likelihood is given by:
$P(y)=\prod_{t=1}^{T}P(y^{(t)}\mid y^{(<t)}),$
where $y^{(<t)}$ denotes the preceding tokens. In practice, we compute the log-likelihood (negative cross-entropy loss) for each sentence and use it as a proxy for model-assigned likelihood (Appendix G).
@@ 5.3 Results
We present the frequency results for GPT-3.5 Turbo, Llama 2 (70B) and Flan-T5 XL in Figure 3. For the likelihood analysis, we present the results for Llama 3 (8B) and Flan-T5 XXL in Figure 4. Results for the other models can be found in Appendix H. These models were chosen as they are representative of the general patterns observed across all the models evaluated.
Frequency is not a free lunch.
In general, we do not observe a consistent pattern that accuracy correlates with expression frequency, suggesting that while models may have encountered certain expressions frequently during pretraining, they may still struggle to properly interpret their idiomaticity in new contexts. Notably, this trend was observed in 9 out of the 13 models we analyzed. Additionally, as the frequency of an expression increases, the models tend to perform better at identifying its literal occurrences, but their accuracy in recognizing idiomatic uses declines. One possible explanation is that for high-frequency expressions, the model may have seen both literal and idiomatic usages during pretraining. However, due to limited contextual understanding, the model may default to interpreting these expressions literally more often, regardless of the actual usage; however, this hypothesis requires further investigation. The contrasting relationship between the two settings explains why no overall correlation is observed between accuracy and frequency. Therefore, frequency does not guarantee performance.
Likelihood $\neq$ Understanding.
Our analysis of sentence likelihoods reveals contrasting trends across model families. In the case of the Llama models, performance correlates positively with sentence probability. In both settings, the model perform better on sentences on which it has a higher likelihood. This is particularly the case on the literal subsets, as indicated by Llama 2 (13B) and Llama 3 (8B). For the Flan-T5 models, we see a negative or negligible correlation between frequency/likelihood and performance. As seen in § 4.2, Small and Large do not appear to have effectively utilized the context to learn meanings as successfully as the other models. The counter-intuitive pattern observed in XL and XXL models could be due to models being over-confident in their wrong prediction. This leads to situations where model would assign high probabilities to idiomatic sentences but performs poorly on the idiomaticity detection task, where surface-level fluency inflates likelihood scores without supporting deeper semantic resolution. These results show that high likelihood does not necessarily imply correct contextual understanding, especially for figurative language.
@@ 6 Conclusion
In this work, we contribute to idiomaticity detection in NLP with several key findings. First, we introduce DICE, a challenging dataset for context-dependent idiom detection, where distinguishing figurative from literal meanings relies heavily on understanding context. Second, we propose an evaluation framework that measures both overall accuracy and strict consistency, requiring models to correctly identify all figurative and literal instances of an expression across different contexts. Third, we show that current LLMs struggle to use context effectively, highlighting the need for models that better capture contextual nuances.
We also investigate the effects of expression frequency and sentence likelihood. While frequency can correlate with performance in some settings, it does not guarantee accurate interpretation, highlighting that surface-level exposure is not a substitute for contextual reasoning. Similarly, while some models tend to perform better on sentences with higher likelihood, this correlation is inconsistent across models. Overall, our findings highlight the limitations of existing models in comprehending idiomatic language and highlight the need for evaluation settings, and model architectures, that emphasize deep contextual understanding over memorization or distributional familiarity.
@@ 7 Limitations
One of the limitations of our work is that some idiomatic expressions are noticeably more reliant on the context than others. This means that there were cases, where we could not provide a literal counterpart to the figurative interpretation. For example, the expression “set eyes on” has such a dominant meaning of “to see”, that the annotators believed to be impossible to override. In these cases, we would discard the expression. As a result, our dataset only contains a selected sample of idioms, and we acknowledge that this idea of contrastive evaluation cannot necessarily be applied to all idioms in a language.
Another of the limitations of our work is that we only consider English idioms. We would like to have extended this work to other languages, however, this relies on the existence of idiomaticity datasets in the target languages. Moreover, the idea of making idioms literal might not be translatable to other languages, where the expression’s domainant, figurative meaning cannot be overridden.
@@ Acknowledgements
We thank the anonymous reviewers and metareviewer for their feedback on this paper. MM is supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1]. AV is supported by UK EPSRC grant EP/T02450X/1, The Alan Turing Institute Fellowship, the UniDive COST Action and MRC-FAPESP MR/U506734/1. We acknowledge IT Services at The University of Sheffield for the provision of services for High Performance Computing. Additional thanks to our annotators for their hard work. A further acknowledgement to Lily Zeng, Thomas Pickard and members of the MWE Group for their helpful discussions.
@@ Appendix A Examples of Expert Annotations
Idiom Definition of the Figurative Meaning Sentence Accept Reject Reason (if reject) smoking gun "a piece of incontrovertible evidence" The detective found a smoking gun at the crime scene. N Y Ambiguous guilt trip "to make someone feel guilty" After breaking her mother’s vase, Sarah’s sister put her on a guilt trip for weeks. N Y Doesn’t make sense turn a blind eye "pretend not to notice" Despite the obvious safety hazards, the supervisor chose to turn a blind eye. N Y Figurative down the wire "a situation whose outcome is not decided until the very last minute" The electrician was careful not to cut down to the wire while he was working. N Y Form changed set eyes on "see" As soon as she set eyes on the beach, she was overwhelmed by its serene beauty. N Y Skip blow off steam "get rid of pent-up energy or emotion" During the train ride, the kids were excited to see the old locomotive blow off steam. Y get a grip "begin to deal with or understand" He struggled to get a grip on the slippery glass jar of pickles. Y
@@ Appendix D Participant Briefing
Appendix E Evaluation Implementation Details
E.1 ChatGPT Versions
We evaluate the following GPT models: GPT-4o (gpt-4o-2024-08-06)  [^2] https://platform.openai.com/docs/models/gpt-4o, GPT-3.5-Turbo (gpt-3_5-turbo-0125) [^3] https://platform.openai.com/docs/models/gpt-3.5-turbo and GPT-4 (gpt-4-0613) [^4] https://platform.openai.com/docs/models/gpt-4.
E.2 Prompting Paradigm
We ran the FLAN-T5 models on a NVIDIA H100 GPU. We use OpenAI’s API for interactions with the GPT models, HuggingFace for Flan-T5 models and Replicate  [^5] https://replicate.com/ for the Llama models. Each model was evaluated with three different prompts. All of the results we report are the average across the three prompt settings.
Method Prompt No. Prompt Design
Zero-shot Prompt 1 Is the expression ’idiom’ used figuratively or literally in the sentence: ’sentence’. Answer ’i’ for figurative, ’l’ for literal.
Prompt 2 In the sentence ’sentence’, is the expression ’idiom’ being used figuratively or literally? Respond with ’i’ for figurative and ’l’ for literal.
Prompt 3 How is the expression ’idiom’ used in this context: ’sentence’. Output ’i’ if the expression holds figurative meaning, output ’l’ if the expression holds literal meaning.
Example The expression ’play with fire’ is used figuratively in the sentence: ’The war took away the unfortunate necessity, as Unionists saw it, to play with fire in the national interest, but it did not materially alter their view of themselves.’ → Output: i The expression ’play with fire’ is used literally in the sentence: ’Despite the danger, he decided to play with fire, poking the embers with a stick.’ → Output: l"
Few-shot Prompt 1 Example + Prompt 1
Prompt 2 Example + Prompt 2
Prompt 3 Example + Prompt 3
Table 5: We use three prompts for our experiments. The top panel of the table shows the prompts used in the zero-shot setting, while the bottom panel displays the few-shot prompts. For the few-shot setting, we prepend the same example (middle panel) to each of the zero-shot prompts.
Figure 6: Left to right: Frequency analysis for Llama 2 (70B), Llama 3 (70B), and Llama 3.1 (405B)
@@ E.3 Hyper-parameters
Table 6: Hyper-parameters used for model evaluation, sorted by model family
@@
Appendix F Frequency Counts
We access the enTenTen corpus using SketchEngine and employ Corpus Query Language (CQL) to find concordances that match specific lexical patterns.
For each target expression, we determine its frequency in the corpus by accounting for all lemma-based forms of the expression. We simply slot in the lemmas of the target expression directly into the following CQL query pattern. For example, for the expression “spill the beans”, the CQL would be: [lemma=‘‘spill’’][lemma=‘‘the’’][lemma=‘‘bean’’]. This would capture occurrences such as “spilled the beans”, “spilling the beans”, and other morphological variants. We utilize NLTK Bird and Loper (2004) to perform lemmatisation and acquire the lemmas needed for the CQL query.
It is important to note that CQL query we use does not account for more syntactically flexible realizations, such as passive constructions (e.g., “the beans were spilled”), as these deviate from the fixed linear ordering captured by the CQL query. Consequently, the resulting frequency estimates represent a conservative measure. Nonetheless, for the purposes of correlating expression frequency with downstream model performance, we consider this approximation to be sufficiently informative.
Appendix G Sentence Likelihood
We derive the sentence likelihood by using the cross-entropy loss. For a sequence of tokens $y=[y^{(1)},y^{(2)},\dots,y^{(T)}]$, the sentence-level cross-entropy loss $L_{\text{sentence}}$ is defined as:
$L_{\text{sentence}}=-\sum_{t=1}^{T}\log P(y^{(t)}\mid y^{(<t)})$ (8)
where $y^{<t}$ represents all tokens preceding the token at position $t$.
Recognizing the relationship between the cross-entropy loss and the sequence probability, we observe:
$L_{\text{sentence}}$ $=-\log\left(\prod_{t=1}^{T}P(y^{(t)}\mid y^{(<t)})\right)$ (9)
$=-\log P(y)$ (10)
Thus, the log likelihood of the sentence is
$\log P(y)=-L_{\text{sentence}}$ (11)
We calculate the log likelihood for all sentences in DICE.
Appendix H Additional Results for Frequency and Likelihood Analysis